OpenStudy (anonymous):

4 − 5x^3 + 5x − 5 = 0

OpenStudy (saifoo.khan):

Find..?

OpenStudy (saifoo.khan):

roots? or factor?

OpenStudy (anonymous):

all solutions

OpenStudy (anonymous):

\[5 (x-1) (x+1) \left(x^2-x+1\right) \]

OpenStudy (anonymous):

so.. if robtobey is correct.. x=1 x=-1 x= (1 +j2)/2 x= (1 -j2)/2